# Scalable Decentralized Learning with Teleportation

### v1

@@ 3.2 Convergence Analysis of Teleportation
$k=\max\left\{1,\min\left\{\Bigg{\lceil}\left(\frac{T(\sigma^{2}+\zeta^{2})}{Lr_{0}}\right)^{\frac{1}{7}}\Bigg{\rceil},\Bigg{\lceil}\left(\frac{T(\sigma^{2}+\zeta^{2})}{Lr_{0}}\right)^{\frac{1}{5}}\Bigg{\rceil},n\right\}\right\},$
@@ B.3 Main Proof
$\leq\mathcal{O}\left(\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{3}{4}}}{T}\right)^{\frac{4}{7}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{2}{3}}}{T}\right)^{\frac{3}{5}}+\frac{Lr_{0}}{T}\right),$
@@
$k=\max\left\{1,\min\left\{\Bigl{\lceil}\left(\frac{A^{3}}{B^{6}}\right)^{\frac{1}{7}}\Bigr{\rceil},\Bigl{\lceil}\left(\frac{A}{C^{2}}\right)^{\frac{1}{5}}\Bigr{\rceil},n\right\}\right\}.$
@@
$k=\min\left\{\Bigl{\lceil}\left(\frac{A^{3}}{B^{6}}\right)^{\frac{1}{7}}\Bigr{\rceil},\Bigl{\lceil}\left(\frac{A}{C^{2}}\right)^{\frac{1}{5}}\Bigr{\rceil},n\right\}.$
@@
    $\sqrt{\frac{A}{k}}=\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}\right),\quad Bk^{\frac{2}{3}}=\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}\right),\quad Ck^{2}$ $\leq\mathcal{O}\left(A^{\frac{2}{5}}C^{\frac{1}{5}}\right),$
    where we use $k\leq\Bigl{\lceil}\left(\frac{A}{C^{2}}\right)^{\frac{1}{5}}\Bigr{\rceil}$. By using the above inequalities, $\tfrac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\|\nabla f(\bar{{\bm{x}}}^{(t)}_{\text{active}})\|^{2}$ is bounded from above by
    $\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}+A^{\frac{2}{5}}C^{\frac{1}{5}}\right).$
  • When $k=\Bigl{\lceil}\left(\frac{A}{C^{2}}\right)^{\frac{1}{5}}\Bigr{\rceil}$, we have
    $\sqrt{\frac{A}{k}}=\mathcal{O}\left(A^{\frac{2}{5}}C^{\frac{1}{5}}\right),\quad Bk^{\frac{2}{3}}\leq\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}\right),\quad Ck^{2}=\mathcal{O}\left(A^{\frac{2}{5}}C^{\frac{1}{5}}\right),$
@@
    $\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}+A^{\frac{2}{5}}C^{\frac{1}{5}}\right).$
  • When $k=n$, we have
    $\sqrt{\frac{Lr_{0}(\sigma^{2}+(1-\frac{k-1}{n-1})\zeta^{2})}{kT}}=\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}},\quad Bk^{\frac{2}{3}}\leq\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}\right),\quad Ck^{2}\leq\mathcal{O}\left(A^{\frac{2}{5}}C^{\frac{1}{5}}\right),$
    where we use $k\leq\Bigl{\lceil}\left(\frac{A^{3}}{B^{6}}\right)^{\frac{1}{7}}\Bigr{\rceil}$ and $k\leq\Bigl{\lceil}\left(\frac{A}{C^{2}}\right)^{\frac{1}{5}}\Bigr{\rceil}$. By using the above inequalities, $\tfrac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\|\nabla f(\bar{{\bm{x}}}^{(t)}_{\text{active}})\|^{2}$ is bounded from above by
    $\mathcal{O}\left(\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}}+A^{\frac{2}{7}}B^{\frac{3}{7}}+A^{\frac{2}{5}}C^{\frac{1}{5}}\right).$
@@ Appendix F Additional Experiment
Figure 6: Test accuracy of Teleportation and Decentralized SGD under heterogeneous networks with $\tau=0$.
@@ Appendix G Experimental Setup
Table 6: Experimental setups for Figs. 4 and 6.

### v2

@@ 3.2 Convergence Analysis of Teleportation
$k=\max\left\{1,\min\left\{\Bigg{\lceil}\left(\frac{T(\sigma^{2}+\zeta^{2})}{Lr_{0}}\right)^{\frac{1}{7}}\Bigg{\rceil},n\right\}\right\},$
@@
$\mathcal{O}\left(\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{3}{4}}}{T}\right)^{\frac{4}{7}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{2}{5}}}{T}\right)^{\frac{5}{7}}+\frac{Lr_{0}}{T}\right),$
@@ 3.4 Convergence Analysis of Teleportation with Efficient Hyperparameter Search
$\mathcal{O}\left(\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{3}{4}}}{T}\right)^{\frac{4}{7}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{2}{5}}}{T}\right)^{\frac{5}{7}}+\frac{Lr_{0}}{T}\right),$
@@ B.3 Main Proof
$\leq\mathcal{O}\left(\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{3}{4}}}{T}\right)^{\frac{4}{7}}+\left(\frac{Lr_{0}(\sigma^{2}+\zeta^{2})^{\frac{2}{5}}}{T}\right)^{\frac{5}{7}}+\frac{Lr_{0}}{T}\right),$
@@
$k=\max\left\{1,\min\left\{\Bigl{\lceil}\left(\frac{A^{3}}{B^{6}}\right)^{\frac{1}{7}}\Bigr{\rceil},n\right\}\right\}.$
@@
$k=\min\left\{\Bigl{\lceil}\left(\frac{A^{3}}{B^{6}}\right)^{\frac{1}{7}}\Bigr{\rceil},n\right\}.$
@@
    $\sqrt{\frac{A}{k}}=\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}\right),\quad Bk^{\frac{2}{3}}=\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}\right),\quad Ck^{2}$ $=\mathcal{O}\left(CA^{\frac{6}{7}}B^{\frac{-12}{7}}\right).$
    By using the above inequalities, $\tfrac{1}{T+1}\sum_{t=0}^{T}\mathbb{E}\|\nabla f(\bar{{\bm{x}}}^{(t)}_{\text{active}})\|^{2}$ is bounded from above by
    $\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}+A^{\frac{6}{7}}B^{\frac{-12}{7}}C\right).$
  • When $k=n$, we have
    $\sqrt{\frac{Lr_{0}(\sigma^{2}+(1-\frac{k-1}{n-1})\zeta^{2})}{kT}}=\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}},\quad Bk^{\frac{2}{3}}\leq\mathcal{O}\left(A^{\frac{2}{7}}B^{\frac{3}{7}}\right),\quad Ck^{2}\leq\mathcal{O}\left(CA^{\frac{6}{7}}B^{\frac{-12}{7}}\right),$
@@
    $\mathcal{O}\left(\sqrt{\frac{Lr_{0}\sigma^{2}}{nT}}+A^{\frac{2}{7}}B^{\frac{3}{7}}+A^{\frac{6}{7}}B^{\frac{-12}{7}}C\right).$
@@ Appendix F Additional Experiment
Figure 6: The behavior of consensus error $\frac{1}{k}\sum_{v_{i}\in V^{(t)}_{\text{active}}}\|{\bm{x}}_{i}^{(t)}-\bar{{\bm{x}}}^{(t)}_{\text{active}}\|^{2}$ for different stochastic noise $\sigma^{2}$ and heterogeneity $\zeta^{2}$. The experimental setup was the same as in Fig. 2. Note that when the optimal number of active nodes $k$ is one, we did not plot the consensus error. See Table 1.
Figure 7: Test accuracy of TELEPORTATION and Decentralized SGD under heterogeneous networks with $\tau=0$. The experimental setup was the same as in Fig. 4.
@@ Appendix G Experimental Setup
Table 6: Experimental setups for Figs. 4 and 7.